A boat travels at a constant rate of 8 miles per hour in still water. It travels upstream for 3/4 of an hour. It then turns around and travels downstream, returning to the starting point, in 1/2 of an hour. Find the rate of the current.
What's the answer and how do you solve it. This is driving me crazy! :-D Please Help

Answer by scott8148(6628)

rate times time equals distance __ the distances are the same

let x=current __ (8-x)(3/4)=(8+x)(1/2)

multiplying by 4 to clear fractions __ 24-3x=16+2x

8=5x